# Representations of a quantum-deformed Lorentz algebra, Clebsch-Gordan map, and Fenchel-Nielsen representation of quantum complex flat connections at level-$\bm{k}$

### v1

@@ 3.2 The Clebsch-Gordan decomposition
The tensor product representation of $U_{{\bf q}}(sl_{2})\otimes U_{\tilde{{\bf q}}}(sl_{2})$ on $\mathcal{H}\otimes\mathcal{H}$ is given by $(\pi_{\lambda_{2},\tilde{\lambda}_{2}}\otimes\pi_{\lambda_{1},\tilde{\lambda}_{1}})\circ\Delta$. In the following, we often use the notation e.g. $K_{1}\equiv K_{\lambda_{1}}\equiv\pi_{\lambda_{1},\tilde{\lambda}_{1}}(K)$ and similar for other generators. We only focus on the represesentation of untilded operator, while the tilded operator can be analyzed in the same way.
@@
$Q_{21}=$ $K_{2}^{-2}Q_{1}+Q_{2}K_{1}^{2}+\left({\bf q}+{\bf q}^{-1}\right)K_{2}^{-2}K_{1}^{2}+K_{2}^{-1}\left({\bf q}^{-1}K_{2}^{2}+{\bf q}K_{2}^{-2}+Q_{2}\right)E_{2}^{-1}E_{1}K_{1}$
$+K_{2}^{-1}E_{2}\left({\bf q}^{-1}K_{1}^{2}+{\bf q}K_{1}^{-2}+Q_{1}\right)E_{1}^{-1}K_{1}.$
@@
    $\mathcal{S}_{2}f(\mu_{2},m_{2}\mid\mu_{1},m_{1})=$ $e^{\frac{2\pi i}{k}\left(\mu_{2}\bm{\nu}_{1}-m_{2}\bm{n}_{1}\right)}f(\mu_{2},m_{2}\mid\mu_{1},m_{1})$
@@
    $t_{21}=\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}},e^{\tilde{\bm{Y}}_{1}-\tilde{\bm{Y}}_{2}+2\tilde{\bm{X}}_{2}}\right),$
@@
    To understand the action of $t_{21}$ on $\mathcal{H}\otimes\mathcal{H}$, we consider the Weil transformation $\mathcal{V}_{2}$ representing the following symplectic transformation:
    $\qquad\left(\begin{array}[]{c}Y_{2}\\ X_{2}\end{array}\right)\mapsto\left(\begin{array}[]{c}Y_{2}-2X_{2}\\ X_{2}\end{array}\right)=\begin{pmatrix}1&-2\\ 0&1\end{pmatrix}\left(\begin{array}[]{c}Y_{2}\\ X_{2}\end{array}\right)=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}\begin{pmatrix}1&0\\ 2&1\end{pmatrix}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\left(\begin{array}[]{c}Y_{2}\\ X_{2}\end{array}\right),$ (3.20)
@@
    $e^{-\bm{Y}_{2}+2\bm{X}_{2}}\mathcal{V}_{2}f=\mathcal{V}_{2}e^{-\bm{Y}_{2}}f,\qquad e^{-\tilde{\bm{Y}}_{2}+2\tilde{\bm{X}}_{2}}\mathcal{V}_{2}f=\mathcal{V}_{2}e^{-\tilde{\bm{Y}}_{2}}f,$
@@
    $t_{21}=\mathcal{V}_{2}\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}},e^{\tilde{\bm{Y}}_{1}-\tilde{\bm{Y}}_{2}}\right)\mathcal{V}_{2}^{-1},$
    where $\varphi(e^{\bm{Y}_{1}-\bm{Y}_{2}},e^{\tilde{\bm{Y}}_{1}-\tilde{\bm{Y}}_{2}})$ simply multiplies the quantum dilogarithm function $\varphi(y_{1}y_{2}^{-1},\tilde{y}_{1}\tilde{y}_{2}^{-1})$ to $f(\mu_{2},m_{2}\mid\mu_{1},m_{1})$.
@@
    $C_{1}^{-1}=\varphi\left(e^{\bm{Y}_{1}-L_{2}},e^{\tilde{\bm{Y}}_{1}-\tilde{L}_{2}}\right)\frac{\varphi\left(e^{-2\bm{X}_{1}+L_{1}},e^{-2\tilde{\bm{X}}_{1}+\tilde{L}_{1}}\right)}{\varphi\left(e^{2\bm{X}_{1}+L_{1}},e^{2\tilde{\bm{X}}_{1}+\tilde{L}_{1}}\right)}$
@@
The unitary transfromation $t_{12}\mathcal{S}_{2}$ transforms $Q_{21}$ to act only on the second factor of $\mathcal{H}\otimes\mathcal{H}$. Namely $Q_{21}=t_{21}\mathcal{S}_{2}Q_{1}^{\prime}\mathcal{S}_{2}^{-1}t_{21}^{-1}$ where
$Q_{1}^{\prime}=e^{\bm{Y}_{1}-2\bm{X}_{1}}+e^{-\bm{Y}_{1}+2\bm{X}_{1}}+e^{-\bm{Y}_{1}}Q_{1}+e^{-2\bm{X}_{1}}Q_{2}+e^{-2\bm{X}_{1}-\bm{Y}_{1}}.$ (3.21)
@@
Since $\mathcal{S}_{2}e^{\bm{Y}_{1}}\mathcal{S}_{2}^{-1}=e^{\bm{Y}_{1}-\bm{Y}_{2}}$ and $\mathcal{S}_{2}e^{\bm{X}_{1}}\mathcal{S}_{2}^{-1}=e^{\bm{X}_{1}}$, we have
$\mathcal{S}_{2}Q_{1}^{\prime}\mathcal{S}_{2}^{-1}=e^{\bm{Y}_{1}-\bm{Y}_{2}-2\bm{X}_{1}}+e^{-\bm{Y}_{1}+\bm{Y}_{2}+2\bm{X}_{1}}+e^{-\bm{Y}_{1}+\bm{Y}_{2}}Q_{1}+e^{-2\bm{X}_{1}}Q_{2}+e^{-2\bm{X}_{1}-\bm{Y}_{1}+\bm{Y}_{2}}.$
Since $[\bm{Y}_{1}-\bm{Y}_{2}-2\bm{X}_{1},\,\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}]=[-\bm{Y}_{2},2\bm{X}_{2}]+[-2\bm{X}_{1},\,\bm{Y}_{1}]=0$, ther first two terms commute with $t_{21}=\varphi(e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}},e^{\tilde{\bm{Y}}_{1}-\tilde{\bm{Y}}_{2}+2\tilde{\bm{X}}_{2}})$. In the following, we often suppress the tilded entry of $\varphi$ when it is not involved in the manipulation. We check the following relation by using the recursion relation of $\varphi(y,\tilde{y})$:
$t_{21}e^{-\bm{Y}_{1}+\bm{Y}_{2}}t_{21}^{-1}$ $=\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)e^{-\bm{Y}_{1}+\bm{Y}_{2}}\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)^{-1}$
$=\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)\varphi\left({\bf q}^{-2}e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)^{-1}e^{-\bm{Y}_{1}+\bm{Y}_{2}}$
$=\left(1+{\bf q}^{-1}e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)e^{-\bm{Y}_{1}+\bm{Y}_{2}}$
$=e^{-\bm{Y}_{1}+\bm{Y}_{2}}+e^{2\bm{X}_{2}}$
$t_{21}e^{-2\bm{X}_{1}}t_{21}^{-1}$ $=\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)e^{-2\bm{X}_{1}}\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)^{-1}$
$=e^{-\bm{X}_{1}}\varphi\left({\bf q}e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)\varphi\left({\bf q}^{-1}e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)^{-1}e^{-\bm{X}_{1}}$
$=e^{-\bm{X}_{1}}\left(1+e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)e^{-\bm{X}_{1}}$
$=e^{-2\bm{X}_{1}}+e^{\bm{Y}_{1}-2\bm{X}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}$
$t_{21}e^{-2\bm{X}_{1}-\bm{Y}_{1}+\bm{Y}_{2}}t_{21}^{-1}$ $=\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)e^{-2\bm{X}_{1}-\bm{Y}_{1}+\bm{Y}_{2}}\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)^{-1}$
$=\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)\varphi\left({\bf q}^{-4}e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)^{-1}e^{-2\bm{X}_{1}-\bm{Y}_{1}+\bm{Y}_{2}}$
$=\left(1+{\bf q}^{-3}e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)\left(1+{\bf q}^{-1}e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)e^{-2\bm{X}_{1}-\bm{Y}_{1}+\bm{Y}_{2}}$
$=\left[1+\left({\bf q}^{-1}+{\bf q}^{-3}\right)e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}+{\bf q}^{-4}e^{2\bm{Y}_{1}-2\bm{Y}_{2}+4\bm{X}_{2}}\right]e^{-2\bm{X}_{1}-\bm{Y}_{1}+\bm{Y}_{2}}$
$=e^{-2\bm{X}_{1}-\bm{Y}_{1}+\bm{Y}_{2}}+\left({\bf q}+{\bf q}^{-1}\right)e^{-2\bm{X}_{1}+2\bm{X}_{2}}+e^{\bm{Y}_{1}-\bm{Y}_{2}-2\bm{X}_{1}+4\bm{X}_{2}}$
@@
$t_{21}\mathcal{S}_{2}Q_{1}^{\prime}\mathcal{S}_{2}^{-1}t_{21}^{-1}=$ $\bm{x}_{0,-1}^{(2)}\bm{x}_{-2,1}^{(1)}+\bm{x}_{0,1}^{(2)}\bm{x}_{2,-1}^{(1)}+Q_{1}\left(\bm{x}_{0,1}^{(2)}\bm{x}_{0,-1}^{(1)}+\bm{x}_{2,0}^{(2)}\right)+Q_{2}\left(\bm{x}_{-2,0}^{(1)}+\bm{x}_{2,-1}^{(2)}\bm{x}_{-2,1}^{(1)}\right)$
$+\bm{x}_{0,1}^{(2)}\bm{x}_{-2,-1}^{(1)}+\left({\bf q}+{\bf q}^{-1}\right)\bm{x}_{2,0}^{(2)}\bm{x}_{-2,0}^{(1)}+\bm{x}_{4,-1}^{(2)}\bm{x}_{-2,1}^{(1)}.$
It is the same as the expression of $Q_{21}$ in terms of $\bm{x}_{\alpha,\beta}$. ∎
@@
${Q}_{1}^{\prime\prime}=C_{1}Q_{1}^{\prime}C_{1}^{-1}=e^{-2\bm{X}_{1}+L_{2}}+e^{2\bm{X}_{1}-L_{2}}+e^{-\bm{Y}_{1}-L_{1}}.$
@@
$C_{1}\left(e^{\bm{Y}_{1}-2\bm{X}_{1}}+e^{-2\bm{X}_{1}}Q_{2}+e^{-2\bm{X}_{1}-\bm{Y}_{1}}\right)C_{1}^{-1}$
$=$ $\frac{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}\varphi\left(e^{\bm{Y}_{1}-L_{2}}\right)^{-1}e^{-\bm{X}_{1}}\left(e^{\bm{Y}_{1}}+e^{-\bm{Y}_{1}}+e^{L_{2}}+e^{-L_{2}}\right)e^{-\bm{X}_{1}}\varphi\left(e^{\bm{Y}_{1}-L_{2}}\right)\frac{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}$
$=$ $e^{-\bm{X}_{1}}\frac{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}\varphi\left({\bf q}e^{\bm{Y}_{1}-L_{2}}\right)^{-1}e^{-\bm{Y}_{1}}\left(1+e^{\bm{Y}_{1}-L_{2}}\right)\left(1+e^{\bm{Y}_{1}+L_{2}}\right)\varphi\left({\bf q}^{-1}e^{\bm{Y}_{1}-L_{2}}\right)\frac{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}e^{-\bm{X}_{1}}$
$=$ $e^{-\bm{X}_{1}}\frac{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}e^{-\bm{Y}_{1}}\left(1+e^{\bm{Y}_{1}+L_{2}}\right)\frac{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}e^{-\bm{X}_{1}}$
$=$ $\frac{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}e^{-\bm{X}_{1}}e^{-\bm{Y}_{1}}e^{-\bm{X}_{1}}\frac{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}+e^{-2\bm{X}_{1}+L_{2}}.$

### v2

@@ 3.2 The Clebsch-Gordan decomposition
The tensor product representation of $\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$ on $\mathcal{H}\otimes\mathcal{H}$ is given by $(\pi_{\lambda_{2},\tilde{\lambda}_{2}}\otimes\pi_{\lambda_{1},\tilde{\lambda}_{1}})\circ\Delta$. In the following, we often use the notation e.g. $K_{1}\equiv K_{\lambda_{1}}\equiv\pi_{\lambda_{1},\tilde{\lambda}_{1}}(K)$ and similar for other generators. We only focus on the represesentation of untilded operator, while the tilded operator can be analyzed in the same way.
@@
$Q_{21}$ $=$ $K_{2}^{-1}Q_{1}+Q_{2}K_{1}+\left(\mathbf{q}+\mathbf{q}^{-1}\right)K_{2}^{-1}K_{1}+\left(Q_{2}+\mathbf{q}^{-1}K_{2}+\mathbf{q}K_{2}^{-1}\right)F_{2}^{-1}K_{2}^{-1}K_{1}F_{1}$ (3.22)
$+\,F_{2}\left(Q_{1}+\mathbf{q}^{-1}K_{1}+\mathbf{q}K_{1}^{-1}\right)F_{1}^{-1}$
@@
    $\mathcal{S}_{2}f(\mu_{2},m_{2}\mid\mu_{1},m_{1})=$ $e^{\frac{2\pi i}{N}\left(\mu_{2}\bm{\nu}_{1}-m_{2}\bm{n}_{1}\right)}f(\mu_{2},m_{2}\mid\mu_{1},m_{1})$
@@
    $t_{21}=\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}},e^{\tilde{\bm{Y}}_{1}-\tilde{\bm{Y}}_{2}+\tilde{\bm{U}}_{2}}\right),$
@@
    To understand the action of $t_{21}$ on $\mathcal{H}\otimes\mathcal{H}$, we consider the Weil transformation $\mathcal{V}_{2}$ representing the following symplectic transformation [12]:
    $\qquad\left(\begin{array}[]{c}Y_{2}\\ U_{2}\end{array}\right)\mapsto\left(\begin{array}[]{c}Y_{2}-U_{2}\\ X_{2}\end{array}\right)=\begin{pmatrix}1&-1\\ 0&1\end{pmatrix}\left(\begin{array}[]{c}Y_{2}\\ U_{2}\end{array}\right)=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}\begin{pmatrix}1&0\\ 1&1\end{pmatrix}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\left(\begin{array}[]{c}Y_{2}\\ X_{2}\end{array}\right),$ (3.31)
@@
    $\mathcal{V}_{2}^{-1}e^{-\bm{Y}_{2}+\bm{U}_{2}}\mathcal{V}_{2}f=-e^{-\bm{Y}_{2}}f,\qquad\mathcal{V}_{2}^{-1}e^{-\tilde{\bm{Y}}_{2}+\tilde{\bm{U}}_{2}}\mathcal{V}_{2}f=-e^{-\tilde{\bm{Y}}_{2}}f,$ (3.32)
@@
    $t_{21}=\mathcal{V}_{2}\varphi\left(-e^{\bm{Y}_{1}-\bm{Y}_{2}},-e^{\tilde{\bm{Y}}_{1}-\tilde{\bm{Y}}_{2}}\right)\mathcal{V}_{2}^{-1},$
    where $\varphi(-e^{\bm{Y}_{1}-\bm{Y}_{2}},-e^{\tilde{\bm{Y}}_{1}-\tilde{\bm{Y}}_{2}})$ simply multiplies the quantum dilogarithm function $\varphi(-y_{1}y_{2}^{-1},-\tilde{y}_{1}\tilde{y}_{2}^{-1})$ to $f(\mu_{2},m_{2}\mid\mu_{1},m_{1})$.
@@
    $C_{1}^{-1}=\varphi\left(e^{\bm{Y}_{1}-L_{2}},e^{\tilde{\bm{Y}}_{1}-\tilde{L}_{2}}\right)\frac{\varphi\left(e^{-\bm{U}_{1}+L_{1}},e^{-\tilde{\bm{U}}_{1}+\tilde{L}_{1}}\right)}{\varphi\left(e^{\bm{U}_{1}+L_{1}},e^{\tilde{\bm{U}}_{1}+\tilde{L}_{1}}\right)}$
@@
The unitary transformation $t_{12}\mathcal{S}_{2}$ transforms $Q_{21}$ to act only on the second factor of $\mathcal{H}\otimes\mathcal{H}$. Namely $Q_{21}=t_{21}\mathcal{S}_{2}Q_{1}^{\prime}\mathcal{S}_{2}^{-1}t_{21}^{-1}$ where
$Q_{1}^{\prime}=e^{\bm{Y}_{1}-\bm{U}_{1}}+e^{-\bm{Y}_{1}+\bm{U}_{1}}+e^{-\bm{Y}_{1}}Q_{1}+e^{-\bm{U}_{1}}Q_{2}+e^{-\bm{U}_{1}-\bm{Y}_{1}}.$ (3.33)
@@
Since $\mathcal{S}_{2}e^{\bm{Y}_{1}}\mathcal{S}_{2}^{-1}=e^{\bm{Y}_{1}-\bm{Y}_{2}}$ and $\mathcal{S}_{2}e^{\bm{U}_{1}}\mathcal{S}_{2}^{-1}=e^{\bm{U}_{1}}$, we have
$\mathcal{S}_{2}Q_{1}^{\prime}\mathcal{S}_{2}^{-1}=e^{\bm{Y}_{1}-\bm{Y}_{2}-\bm{U}_{1}}+e^{-\bm{Y}_{1}+\bm{Y}_{2}+\bm{U}_{1}}+e^{-\bm{Y}_{1}+\bm{Y}_{2}}Q_{1}+e^{-\bm{U}_{1}}Q_{2}+e^{-\bm{U}_{1}-\bm{Y}_{1}+\bm{Y}_{2}}.$
Since $[\bm{Y}_{1}-\bm{Y}_{2}-\bm{U}_{1},\,\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}]=[-\bm{Y}_{2},\bm{U}_{2}]+[-\bm{U}_{1},\,\bm{Y}_{1}]=0$, the first two terms commute with $t_{21}=\varphi(e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}},e^{\tilde{\bm{Y}}_{1}-\tilde{\bm{Y}}_{2}+\tilde{\bm{U}}_{2}})$. In the following, we often suppress the tilded entry of $\varphi$ when it is not involved in the manipulation. We check the following relation by using the recursion relation of the quantum dilogarithm $\varphi$:
$t_{21}e^{-\bm{Y}_{1}+\bm{Y}_{2}}t_{21}^{-1}$ $=\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)e^{-\bm{Y}_{1}+\bm{Y}_{2}}\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)^{-1}$
$=\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)\varphi\left({\bf q}^{-2}e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)^{-1}e^{-\bm{Y}_{1}+\bm{Y}_{2}}$
$=\left(1+{\bf q}^{-1}e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)e^{-\bm{Y}_{1}+\bm{Y}_{2}}$
$=e^{-\bm{Y}_{1}+\bm{Y}_{2}}+e^{\bm{U}_{2}},$
$t_{21}e^{-\bm{U}_{1}}t_{21}^{-1}$ $=\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)e^{-\bm{U}_{1}}\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)^{-1}$
$=\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)\varphi\left({\bf q}^{-2}e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)^{-1}e^{-\bm{U}_{1}}$
$=\left(1+{\bf q}^{-1}e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)e^{-\bm{U}_{1}}$
$=e^{-\bm{U}_{1}}+e^{\bm{Y}_{1}-\bm{U}_{1}-\bm{Y}_{2}+\bm{U}_{2}},$
$t_{21}e^{-\bm{U}_{1}-\bm{Y}_{1}+\bm{Y}_{2}}t_{21}^{-1}$ $=\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)e^{-\bm{U}_{1}-\bm{Y}_{1}+\bm{Y}_{2}}\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)^{-1}$
$=\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)\varphi\left({\bf q}^{-4}e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)^{-1}e^{-\bm{U}_{1}-\bm{Y}_{1}+\bm{Y}_{2}}$
$=\left(1+{\bf q}^{-3}e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)\left(1+{\bf q}^{-1}e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)e^{-\bm{U}_{1}-\bm{Y}_{1}+\bm{Y}_{2}}$
$=\left[1+\left({\bf q}^{-1}+{\bf q}^{-3}\right)e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}+{\bf q}^{-4}e^{2\bm{Y}_{1}-2\bm{Y}_{2}+2\bm{U}_{2}}\right]e^{-\bm{U}_{1}-\bm{Y}_{1}+\bm{Y}_{2}}$
$=e^{-\bm{U}_{1}-\bm{Y}_{1}+\bm{Y}_{2}}+\left({\bf q}+{\bf q}^{-1}\right)e^{-\bm{U}_{1}+\bm{U}_{2}}+e^{\bm{Y}_{1}-\bm{Y}_{2}-\bm{U}_{1}+2\bm{U}_{2}}.$
@@
$t_{21}\mathcal{S}_{2}Q_{1}^{\prime}\mathcal{S}_{2}^{-1}t_{21}^{-1}$ $=$ $\bm{u}_{0,-1}^{(2)}\bm{u}_{-1,1}^{(1)}+\bm{u}_{0,1}^{(2)}\bm{u}_{1,-1}^{(1)}+Q_{1}\left(\bm{u}_{0,1}^{(2)}\bm{u}_{0,-1}^{(1)}+\bm{u}_{1,0}^{(2)}\right)+Q_{2}\left(\bm{u}_{-1,0}^{(1)}+\bm{u}_{1,-1}^{(2)}\bm{u}_{-1,1}^{(1)}\right)$ (3.34)
$+\,\bm{u}_{0,1}^{(2)}\bm{u}_{-1,-1}^{(1)}+\left({\bf q}+{\bf q}^{-1}\right)\bm{u}_{1,0}^{(2)}\bm{u}_{-1,0}^{(1)}+\bm{u}_{2,-1}^{(2)}\bm{u}_{-1,1}^{(1)}.$
It is the same as the expression of $Q_{21}$ in terms of $\bm{u}_{\alpha,\beta}$. ∎
@@
${Q}_{1}^{\prime\prime}=C_{1}Q_{1}^{\prime}C_{1}^{-1}=e^{-\bm{U}_{1}+L_{2}}+e^{\bm{U}_{1}-L_{2}}+e^{-\bm{Y}_{1}-L_{1}}.$
@@
$C_{1}\left(e^{\bm{Y}_{1}-\bm{U}_{1}}+e^{-\bm{U}_{1}}Q_{2}+e^{-\bm{Y}_{1}-\bm{U}_{1}}\right)C_{1}^{-1}$
$=$ $\frac{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}\varphi\left(e^{\bm{Y}_{1}-L_{2}}\right)^{-1}{\bf q}e^{-\bm{Y}_{1}}\left(1+{\bf q}^{-1}e^{\bm{Y}_{1}-L_{2}}\right)\left(1+{\bf q}^{-1}e^{\bm{Y}_{1}+L_{2}}\right)\varphi\left({\bf q}^{-2}e^{\bm{Y}_{1}-L_{2}}\right)\frac{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}e^{-\bm{U}_{1}}$
$=$ $\frac{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}\varphi\left(e^{\bm{Y}_{1}-L_{2}}\right)^{-1}{\bf q}e^{-\bm{Y}_{1}}\left(1+{\bf q}^{-1}e^{\bm{Y}_{1}+L_{2}}\right)\varphi\left(e^{\bm{Y}_{1}-L_{2}}\right)\frac{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}e^{-\bm{U}_{1}}$
$=$ $\frac{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}\left({\bf q}^{-1}e^{-\bm{U}_{1}}e^{-\bm{Y}_{1}}\right)\frac{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}+e^{-\bm{U}_{1}+L_{2}}.$